# How Far Are We From AGI?

### v1

@@ Acknowledgement
6.5 Further Considerations during AGI Development
@@
In this section, we present some hopefully thought-provoking questions to encourage further engagement and discussion on how to responsibly approach AGI. Even though there might or will not be any answer to these questions, we nonetheless give some interpretations and ideas for the sake of sharing our own insights about how overcoming these putative limitations can possibly help get us closer to AGI.
@@
How Far Do Researchers Think We Are From AGI? Despite extensive discussion on many facets of AGI, we haven’t yet touched the question of when we and other researchers think it will actually be achieved. Figure 9 shows the poll results from researchers on their thoughts about it at the ICLR 2024 “How Far Are We From AGI” workshop. Even though almost everyone is optimistic about the ultimate arrival of AGI, opinions on the exact time it takes to do so differ quite a lot, which also implies different bottlenecks people are considering. Those who think an extra one or two years would be sufficient might feel that we’ve reached the point where current AI systems are already capable enough and what is left might just be some incremental improvements on the completeness. Those who believe that more than two decades are needed might either feel skeptical about the current general approach to AI or think we still lack fundamental advancement or understanding of intelligence. The discrepancy of people’s opinions on when AGI will be realized is also one of the motivations for why we write this paper in the first place: situating researchers and engineers on the same common ground for contemplating and discussing the vision and possibility of AGI, which we hope will give a much more unified perspective.
@@
Figure 9: Polls Results of Researchers’ Opinions on When AGI Will be Achieved. Among the attending researchers at the ICLR 2024 “How Far Are We From AGI” workshop, a survey is conducted to gather their opinions on how far they think AGI will be achieved. A total of 138 responses are received as above. Interestingly, 37% of researchers think it will take more than 20 years from now on to realize AGI.
@@
Is Autoregressive Generation the Way to AGI? Next-token prediction is the core of most successful large foundation models (Qi et al., 2024; Wu et al., 2024c). This raises the question: can next-token prediction lead to AGI? Essentially, autoregressive generation that utilizes extensive self-supervised data represents a form of massively multi-task learning. By predicting the subsequent word in a given text from the corpus, it addresses tasks ranging from traditional NLP tasks like grammar, lexical semantics, and translation to commonsense reasoning and knowledge-grounded reasoning. Learning input-output relationships, or in-context learning, can be cast as next-word prediction. The relationships in the world are often encoded in words, visual tokens, spacetime patches, or other types of tokens, allowing them to be learned through next-token prediction. The critical query remains: does the spectrum of world knowledge, including implicit knowledge such as intuition, emotion, culture, and artistic expression, get encoded in simplified tokens? Can the auto-regressive approach learn all the causation in addition to correlations within world knowledge? This subject remains a point of ongoing debate.
@@
What are the Biggest Remaining Hurdles?151515 These questions and contents are inspired and summarized from the panel and audience discussions of the ICLR 2024 “How far are we from AGI” workshop. There are several key research questions and hurdles that need to be addressed to make AGI a reality. We summarize the key points made in the panel discussions:
@@
Open-source in the Era of AGI? 15 The ultimate goal of open-sourcing AI systems is to facilitate AI research and improve human well-being. However, as the technology continues growing, there is an urgent need to balance the pros and cons of open-sourcing powerful AI systems, as “it will be too late to do so once we reach AGI”, and hence the big questions are when to decide and who to decide. We also list a couple of interesting opinions regarding open-sourcing AI systems:
  • Analogy of AI open-sourcing to weapon regulation. The accessibility comes with the danger of misuse, just like the legitimization of weapons in some Western countries. It becomes a lot easier to weaken and even remove the safeguarding of an AI system once its full details get exposed to the public, hence making them easily exploitable. An example of such potential misconduct is the development of a highly persuasive AI agent on social media. With the current technology and heavy fine-tuning of specialized internet data, it is not hard to create some system that can maneuver the audience with persuasion. Therefore, open-sourcing needs to consider the vast potential of different fine-tuning and adoption. However, it is worth mentioning that the level of exploitation is very different, depending on many factors and considerations, one of which is the amount of compute resources available for such fine-tuning [^16] Marginal risk replaces absolute risk to account for each individual’s difference. . Therefore, assessing the potential consequences of open-sourcing a specific piece of software is a challenging problem.
@@
  • Incentives for open-sourcing. Making intellectual works public and sharing greater details needs motivation. For a healthy community to promote beneficial open-sourcing, building an incentive mechanism is crucially important. For many individuals, the incentives are much clearer: young researchers and students, for example, consider open-sourcing as a form of ownership claiming and advertisement of their own works. However, the incentives of many other entities like tech companies are much more obscure, often involving technology barriers and monetization. Overall, motivating people to contribute to the open-source community is a collaborative effort, a necessary step towards open AGI.
@@
Future Research Methodology for AGI.15 Researchers should consider various perspectives and approaches to contribute to the future development of powerful and safe AI systems. Instead of focusing solely on one direction, researchers should explore multiple paths to building AGI in order to avoid the “local minimum”. In this sense, different researchers may have different perspectives. In terms of safety, it is crucial to understand the limitations of different approaches according to Yoshua Bengio. Meanwhile, Andrew G. Wilson calls for stepping back to conduct theoretical as well as empirical analysis to gain fundamental insights about the models. From Yejin Choi’s perspective, it is important to gain a precise understanding of language models as well as benchmarks on the road to AGI. Song Han emphasizes the directions of efficient hardware systems and network architectures to boost efficient training and inference of current AI systems. By adopting a diverse range of research approaches, researchers can contribute meaningfully to the advancement of AI.
@@ 7.1 AI for Science Discovery and Research
Moreover, AI has been instrumental in advancing personalized medicine and disease diagnosis. Deep learning models, such as DeepSEA (Zhou and Troyanskaya, 2015), have shown remarkable accuracy in predicting the impact of genetic variations on disease risk, paving the way for targeted interventions. Additionally, models like MedAgents (Tang et al., 2024) have demonstrated the ability to generate personalized treatment recommendations based on multi-agent collaboration to enhance their reasoning.
@@
In addition to these developments, the PACT model (Han et al., 2022) showcases the ability to generate human-readable proofs for complex mathematical statements, paving the way for more accessible and understandable mathematical reasoning. Moreover, the MathQA model (Amini et al., 2019) has been developed to answer open-ended mathematical questions, showcasing the ability of LLMs to engage in mathematical dialogue and provide explanations to complex concepts. This opens up new possibilities for personalized mathematical education and interactive learning experiences.
@@ 7.2 Generative Visual Intelligence
Image Generation Diffusion models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Song et al., 2020; Ho et al., 2020) have emerged as the new state-of-the-art family of deep generative models (Yang et al., 2023f), outperforming generative adversarial networks (GANs) (Goodfellow et al., 2020) which had previously been dominant in the challenging task of image synthesis. Diffusion models learn to generate data by reversing a diffusion process, which gradually adds noise to the data until it reaches a distribution of pure noise. This process is characterized by learning the reverse diffusion steps, effectively denoising the data, through a parameterized model that is trained to estimate the gradient of the log probability of the clean data given the noisy data. At inference time, these models generate new samples by iteratively applying the learned reverse diffusion process, starting from noise and progressively denoising it to produce samples that resemble the data distribution the model was trained on. Notable improvements to diffusion models include reformulating diffusion models to predict noise instead of pixels (Song and Ermon, 2019), introducing classifier-free guidance (Ho and Salimans, 2022), applying diffusion models in the latent space of pre-trained autoencoder (Rombach et al., 2022), and replacing U-Net with transformer-based backbones (Peebles and Xie, 2023). In addition to diffusion models, the Large Vision Model (LVM) (Bai et al., 2023) and Visual AutoRegressive modeling (VAR)  (Tian et al., 2024a) provide auto-regressive learning paradigm based on different image scales that facilitates effective and high-quality image generation.
Video Generation Video diffusion models (Ho et al., 2022) introduce a conditional sampling technique for spatial and temporal video extension. Sora (Brooks et al., 2024) can generate up to a minute of high-fidelity video by training text-conditional diffusion models jointly on videos and images with varying durations, resolutions, and aspect ratios. It compresses videos into a lower-dimensional latent space and decomposes the representation into spacetime patches. Then, given input noisy patches and conditioning information like text prompts, diffusion transformers (Peebles and Xie, 2023) are trained to reconstruct the original, clean patches. Demonstrating the ability to produce videos with 3D consistency, extensive coherence, and object permanence, Sora illustrates the potential of generative models as highly capable simulators of the physical and digital worlds.
@@ 7.4 Decentralized AI
The advancement of hardware accelerators pushes the success of multi-billion or even trillion-scale language models to its pinnacle. Most of the SoTA LLMs currently are trained and served in data centers with 1) high-end infrastructures such as homogeneous accelerators, 2) optimized network topology for super fast interconnection, 3) stable and efficient power supply, and 4) careful maintenance from human experts. However, training a model like GPT-3 (Brown et al., 2020) from scratch still costs way more than what individuals can afford: i.e. full pretraining of a GPT-3 model, which is no longer the most powerful model, is estimated to still take at least months with a thousand V100 GPUs (Lambda, 2023). Serving models also face many challenges when we scale the batch size up without hurting response latency. Moving towards the era of AGI, we need new technology to help overcome the limitations of the current dominant form of model training and serving, one prominent direction of which is to transcend from data center to decentralized AI.
@@
Mitigating the Hardware Constraints One desired property for edge servers is the ability to serve LLMs even with a commodity accelerator. FlexGen (Sheng et al., 2023b) first shows that it is possible to run text generation of large models like OPT-175B on a single 16GB GPU. FlexGen adaptively offloads to aggregate memory and computation from the GPU, CPU, and disk. With efficient patterns searched via linear programming and weight and cache quantization, it can decode OPT-175B at 1 token/s speed with a batch size of 144 with negligible accuracy loss. To maximize the potential of different hardware, MLC-LLM (MLC team, 2023) provides a universal solution that allows any language model to be deployed natively on a diverse set of hardware backends and native applications. For example, MLCChat, an iOS app, can serve some of the latest iPhone and iPad models; a similar APK is also available for Androids (spanning manufacturers like Samsung, Redmi, and Google). The possibility continues to Mac, PC, Linux, and web browsers. Finally, on the hardware side, more and more powerful yet economical chips are developed to face the excitement of edge LLMs, examples including Apple’s M3 series and Qualcomm’s Cloud AI 100 Ultra (supporting 100-billion-parameter models on a single 150-watt card). Last but not least, nuclear batteries (Prelas et al., 2014) have shown their potential to revolutionize the power structure of mobile compute platforms, with a notable claimed battery duration of 50 years without charging (The Economic Times, 2024), which could potentially make edge devices more accessible, stable, and suitable for the diverse applications of LLMs.
@@ 7.5 AI for Coding
The ability to write programs stands as one of the defining hallmarks of AGI. Writing complicated programs shows the skill of an AI system in abstract reasoning and adaptability in addressing diverse tasks. As Alan Turing once pointed out in his seminal work (Turing, 1950), being able to write codes fundamentally indicates that an AI system can exhibit intelligent behavior akin to human cognition, where the manipulation of symbols (following a specific language grammar to implement algorithms) leads to the manifestation of complex thought processes. Hence developing code LLMs for both understanding and generation is crucially important both practically and conceptually for stepping towards AGI.
Code Foundation Models While many models for code generation are pre-trained mostly on code corpus (Allal et al., 2023; Li et al., 2022b; Fried et al., 2023; Li et al., 2023a), more general purpose LLMs that are continually pre-trained or fine-tuned on code become more powerful and capable such as Codex (Chen et al., 2021b), GPT-4 (OpenAI, 2023a), PaLM-Coder (Chowdhery et al., 2022), CodeLlama (Rozière et al., 2024), and also smaller scale models like Phi (Gunasekar et al., 2023). The transition from code-specialized to code-understanding models also indicates that coding is a fundamental skill for AGI, just like many other forms of general knowledge. Beyond code generation, these models are also capable of multi-language reasoning (OpenAI, 2023a; Rozière et al., 2024) and infilling with before and after context (Fried et al., 2023; Bavarian et al., 2022). Code models open up many applications as the programs directly serve as the most efficient machine language to communicate with other systems, which we will discuss in the next section.
However, it is worth mentioning that code evaluation is challenging than pure texts for many reasons:
@@
Code LLM Applications Code foundation models (Chen et al., 2021b; OpenAI, 2023a; Rozière et al., 2024; Chowdhery et al., 2022) have already been extremely capable of conducting many basic code maneuvering such as completion, revision, doc-string generation, commenting, bug finding (Tian et al., 2024b), and code translation (Murali et al., 2024). There are, however, far more exciting applications of these models with no or minimal fine-tuning, which unfolds the possibility of turning many systems into an amalgamation.
@@
  • Software engineering. Many code applications centers around software engineering and AI-assisted coding beyond the basic abilities described above. SWE-Bench (Jimenez et al., 2023) attempts to assess a model’s capability to resolve GitHub issues, a core activity in a rich and sustainable real-world software community. Doing so requires a coordinated understanding of the problem description, the execution environment, comments, and the codebase which often has cross-file dependencies and extremely long contexts. The fact that their fine-tuned SWE-Llama can only resolve the simplest issues highly motivates more complicated and capable code models that can greatly help the software ecosystem. Software safety and reliability have always been the most pivotal questions for engineers: RLSQM (Steenhoek et al., 2023) studies using reinforcement learning with static quality metrics as rewards for training a code LLM that can effectively generate unit tests for a codebase with little test smells while adhering to better practices; besides algorithmic bugs,  (Ullah et al., 2023) gives a comprehensive LLM evaluation on identifying security-related bugs, the result of which suggests that even the most capable LLMs like GPT-4 (OpenAI, 2023a) and Palm-2 (Anil et al., 2023) are still prone to non-deterministic responses, incorrect and unfaithful reasoning, and significant non-robustness. LLMs are also explored for lower-level code optimization and refinement. (Cummins et al., 2023) shows that it is possible to fine-tune Llama to optimize LLVM assembly via generating a set of compiler options, which leads to an optimized program and, at the same time, predicts the instruction counts for fine controls. (Wong et al., 2023) investigates the feasibility of utilizing LLMs for de-compiling (reverse-engineer) a C executable into re-compilable C source codes which are expected to exhibit the same functionality, a process that is extremely tedious, time-consuming, and often requires great expertise. Towards a holistic AI development companion, Github Copilot [^17] https://github.com/features/copilot provides personalized and natural language-based coding assistance to developers spanning all levels of expertise and has been integrated into major development workflows. Cognition AI recently announced Devin [^18] https://www.cognition-labs.com/blog as the first AI software engineer. Equipped with its own command line, code editor, and browser, Devin not only achieves the SoTA performance on SWE-bench (Jimenez et al., 2023) but, more impressively, shows its incredible potential in 1) utilizing unfamiliar technology (e.g. running ControlNet on Modal to produce images based on a blog post), 2) building apps end-to-end (e.g. create the Game of Life on a website deployed to Netlify), 3) setting up codes for train and fine-tune LLMs, 4) addressing bugs and feature requests in open source repositories, and so on. Dakhel et al. (2023) suggest examining the capabilities of AI-assisted programming tools in a more controlled setting where the correctness, efficiency, and similarity to human-written solutions are considered extensively.

### v2

@@ Acknowledgement
6.5 Alternative Perspectives on the AGI Roadmap
@@
In this section, we pose thought-provoking questions to inspire deeper reflection and discussion on responsibly advancing AGI beyond the scope of LLMs. Even though there might or will not be any answer to these questions, we nonetheless give some interpretations and ideas for the sake of sharing our own insights about how overcoming these putative limitations can possibly help get us closer to AGI.
@@
How Far Do Researchers Think We Are From AGI? Despite extensive discussion on many facets of AGI, we haven’t yet touched the question of when we and other researchers think it will actually be achieved. Figure 9 shows the poll results from researchers on their thoughts about it at the ICLR 2024 “How Far Are We From AGI” workshop [^15] https://agiworkshop.github.io. Even though almost everyone is optimistic about the ultimate arrival of AGI, opinions on the exact time it takes to do so differ quite a lot, which also implies different bottlenecks people are considering. Those who think an extra one or two years would be sufficient might feel that we’ve reached the point where current AI systems are already capable enough and what is left might just be some incremental improvements on the completeness. Those who believe that more than two decades are needed might either feel skeptical about the current general approach to AI or think we still lack fundamental advancement or understanding of intelligence. The discrepancy of people’s opinions on when AGI will be realized is also one of the motivations for why we write this paper in the first place: situating researchers and engineers on the same common ground for contemplating and discussing the vision and possibility of AGI, which we hope will give a much more unified perspective.
@@
Figure 9: Polls Results of Researchers’ Opinions on When AGI Will be Achieved. Among the attending researchers at the ICLR 2024 “How Far Are We From AGI” workshop, a survey was conducted to gather their opinions on how far they think AGI will be achieved. A total of 138 responses are received as above. Interestingly, 37% of researchers think it will take more than 20 years from now on to realize AGI.
@@
Is Autoregressive Generation the Way to AGI? Next-token prediction is the core of most successful large foundation models (Qi et al., 2024; Wu et al., 2024c). This raises the question: can next-token prediction lead to AGI? Essentially, autoregressive generation that utilizes extensive self-supervised data represents a form of massively multi-task learning. By predicting the subsequent word in a given text from the corpus, it addresses tasks ranging from traditional NLP tasks like grammar, lexical semantics, and translation to commonsense reasoning and knowledge-grounded reasoning. Learning input-output relationships, or in-context learning, can be cast as next-word prediction. The relationships in the world are often encoded in words, visual tokens, spacetime patches, or other types of tokens, allowing them to be learned through next-token prediction. The critical query remains: does the spectrum of world knowledge, including implicit knowledge such as intuition, emotion, culture, and artistic expression, get encoded in simplified tokens? Can the auto-regressive approach learn all the causation in addition to correlations within world knowledge? Additionally, the popularity of the diffusion model (Gat et al., 2024) poses challenges to the future of autoregressive generation. This type of method does not rely on previously generated data points during the generation process but rather relies on the process of gradually reducing noise to recover the data. The remarkable effect of the diffusion model in generation tasks has also led to its widespread use in real-world applications (Yang et al., 2023f; Chen et al., 2024b). All of this makes whether the autoregressive generation is the way to AGI an ongoing debate.
@@
Figure 10: AI has Gradually Surpassed Humans. We estimate the (cumulative) percentage of human activities in which AI has surpassed humans in terms of competence and efficiency as we start from early embryonic systems around the 1970s to more advanced ones developed recently. At each node along the polyline, we choose one representative work that revolutionized the field, bringing substantial improvements to AI technology. The trend of AI popularity and generality is increasing at a fast rate and we can expect that starting from 2023, with the advent of work such as GPT-4 and Sora, the speed at which AI surpasses humans will increase at an unprecedented speed. This figure serves as an alternative perspective on the AGI Roadmap, which may inspire discussions about the speed of AI development.
Is AGI within closer reach than ever?
As is summarized in Figure 10, the rapid development of AI has enabled its capabilities to surpass human activities in increasingly more fields in our estimation, which indicates that the realization of AGI is getting closer. Hence, it is of great practical significance to revisit the question of how far we are from AGI and how can we responsibly achieve AGI by conducting a comprehensive survey that clearly establishes the expectation of future AGI and elaborates on the gap from our current AI development.
@@
What are the Biggest Remaining Hurdles? There are several key research questions and hurdles that need to be addressed to make AGI a reality.
@@
Open-source in the Era of AGI? The ultimate goal of open-sourcing AI systems is to facilitate AI research and improve human well-being. However, as the technology continues growing, there is an urgent need to balance the pros and cons of open-sourcing powerful AI systems, as “it will be too late to do so once we reach AGI”, and hence the big questions are when to decide and who to decide. We also list a couple of interesting opinions regarding open-sourcing AI systems:
  • Analogy of AI open-sourcing to weapon regulation. The accessibility comes with the danger of misuse, just like the legitimization of weapons in some Western countries. It becomes a lot easier to weaken and even remove the safeguarding of an AI system once its full details get exposed to the public, hence making them easily exploitable. An example of such potential misconduct is the development of a highly persuasive AI agent on social media. With the current technology and heavy fine-tuning of specialized internet data, it is not hard to create some system that can maneuver the audience with persuasion. Therefore, open sourcing needs to consider the vast potential of different fine-tuning and adoption. However, it is worth mentioning that the level of exploitation is very different, depending on many factors and considerations, one of which is the amount of compute resources available for such fine-tuning [^16] Marginal risk replaces absolute risk to account for each individual’s difference. . Therefore, assessing the potential consequences of open-sourcing a specific piece of software is a challenging problem.
@@
  • Incentives for open-sourcing. Making intellectual works public and sharing greater details needs motivation. For a healthy community to promote beneficial open-sourcing, building an incentive mechanism is crucially important. For many individuals, the incentives are much clearer: young researchers and students, for example, consider opensourcing as a form of ownership claiming and advertisement of their own works. However, the incentives of many other entities like tech companies are much more obscure, often involving technology barriers and monetization. Overall, motivating people to contribute to the open-source community is a collaborative effort, a necessary step towards open AGI.
@@
Future Research Methodology for AGI. Researchers should consider various perspectives and approaches to contribute to the future development of powerful and safe AI systems. Instead of focusing solely on one direction, researchers should explore multiple paths to building AGI in order to avoid the “local minimum”. In this sense, different researchers may have different perspectives. In terms of safety, it is crucial to understand the limitations of different approaches according to Yoshua Bengio. Meanwhile, Andrew G. Wilson calls for stepping back to conduct theoretical as well as empirical analysis to gain fundamental insights about the models. From Yejin Choi’s perspective, it is important to gain a precise understanding of language models as well as benchmarks on the road to AGI. Song Han emphasizes the directions of efficient hardware systems and network architectures to boost efficient training and inference of current AI systems. By adopting a diverse range of research approaches, researchers can contribute meaningfully to the advancement of AI.
@@ 7.1 AI for Science Discovery and Research
Moreover, AI has been instrumental in advancing personalized medicine and disease diagnosis. Deep learning models, such as DeepSEA (Zhou and Troyanskaya, 2015), have shown remarkable accuracy in predicting the impact of genetic variations on disease risk, paving the way for targeted interventions. Additionally, models like MedAgents (Tang et al., 2024b) have demonstrated the ability to generate personalized treatment recommendations based on multi-agent collaboration to enhance their reasoning.
@@
In addition to these developments, the PACT model (Han et al., 2022) showcases the ability to generate human-readable proofs for complex mathematical statements, paving the way for more accessible and understandable mathematical reasoning. Moreover, the MathQA model (Amini et al., 2019) has been developed to answer open-ended mathematical questions, showcasing the ability of LLMs to engage in mathematical dialogue and provide explanations for complex concepts. This opens up new possibilities for personalized mathematical education and interactive learning experiences.
@@ 7.2 Generative Visual Intelligence
Image Generation Diffusion models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Song et al., 2020; Ho et al., 2020) have emerged as the new state-of-the-art family of deep generative models (Yang et al., 2023f), outperforming generative adversarial networks (GANs) (Goodfellow et al., 2020) which had previously been dominant in the challenging task of image synthesis. Diffusion models learn to generate data by reversing a diffusion process, which gradually adds noise to the data until it reaches a distribution of pure noise. This process is characterized by learning the reverse diffusion steps, effectively denoising the data, through a parameterized model that is trained to estimate the gradient of the log probability of the clean data given the noisy data. At inference time, these models generate new samples by iteratively applying the learned reverse diffusion process, starting from noise and progressively denoising it to produce samples that resemble the data distribution the model was trained on. Notable improvements to diffusion models include reformulating diffusion models to predict noise instead of pixels (Song and Ermon, 2019), introducing classifier-free guidance (Ho and Salimans, 2022), applying diffusion models in the latent space of pre-trained autoencoder (Rombach et al., 2022), and replacing U-Net with transformer-based backbones (Peebles and Xie, 2023; Jin and Xie, 2024). In addition to diffusion models, the Large Vision Model (LVM) (Bai et al., 2023) and Visual AutoRegressive modeling (VAR)  (Tian et al., 2024a) provide auto-regressive learning paradigm based on different image scales that facilitates effective and high-quality image generation.
Video Generation Video diffusion models (Ho et al., 2022; Xing et al., 2023) introduce a conditional sampling technique for spatial and temporal video extension. Sora (Brooks et al., 2024) can generate up to a minute of high-fidelity video by training text-conditional diffusion models jointly on videos and images with varying durations, resolutions, and aspect ratios. It compresses videos into a lower-dimensional latent space and decomposes the representation into spacetime patches. Then, given input noisy patches and conditioning information like text prompts, diffusion transformers (Peebles and Xie, 2023) are trained to reconstruct the original, clean patches. Demonstrating the ability to produce videos with 3D consistency, extensive coherence, and object permanence, Sora illustrates the potential of generative models as highly capable simulators of the physical and digital worlds.
@@ 7.4 Decentralized AI
The advancement of hardware accelerators pushes the success of multi-billion or even trillion-scale language models to its pinnacle. Most of the SoTA LLMs currently are trained and served in data centers with 1) high-end infrastructures such as homogeneous accelerators, 2) optimized network topology for super fast interconnection, 3) stable and efficient power supply, and 4) careful maintenance from human experts. However, training a model like GPT-3 (Brown et al., 2020) from scratch still costs way more than what individuals can afford: i.e. full pretraining of a GPT-3 model, which is no longer the most powerful model, is estimated to still take at least months with a thousand V100 GPUs (Lambda, 2023). Serving models also face many challenges when we scale the batch size up without hurting response latency. Moving towards the era of AGI, we need new technology to help overcome the limitations of the current dominant form of model training and serving, one prominent direction of which is to transcend from data centers to decentralized AI.
@@
Mitigating the Hardware Constraints One desired property for edge servers is the ability to serve LLMs even with a commodity accelerator. FlexGen (Sheng et al., 2023b) first shows that it is possible to run text generation of large models like OPT-175B on a single 16GB GPU. FlexGen adaptively offloads to aggregate memory and computation from the GPU, CPU, and disk. With efficient patterns searched via linear programming and weight and cache quantization, it can decode OPT-175B at 1 token/s speed with a batch size of 144 with negligible accuracy loss. To maximize the potential of different hardware, MLC-LLM (MLC team, 2023) provides a universal solution that allows any language model to be deployed natively on a diverse set of hardware backends and native applications. For example, MLCChat, an iOS app, can serve some of the latest iPhone and iPad models; a similar APK is also available for Androids (spanning manufacturers like Samsung, Redmi, and Google). The possibility continues to Mac, PC, Linux, and web browsers. Finally, on the hardware side, more and more powerful yet economical chips are developed to face the excitement of edge LLMs, examples including Apple’s M3 series and Qualcomm’s Cloud AI 100 Ultra (supporting 100-billion-parameter models on a single 150-watt card). Last but not least, nuclear batteries (Prelas et al., 2014) have shown their potential to revolutionize the power structure of mobile computing platforms, with a notable claimed battery duration of 50 years without charging (The Economic Times, 2024), which could potentially make edge devices more accessible, stable, and suitable for the diverse applications of LLMs.
@@ 7.5 AI for Coding
The ability to write programs stands as one of the defining hallmarks of AGI. Writing complicated programs shows the skill of an AI system in abstract reasoning and adaptability in addressing diverse tasks. As Alan Turing once pointed out in his seminal work (Turing, 1950), being able to write codes fundamentally indicates that an AI system can exhibit intelligent behavior akin to human cognition, where the manipulation of symbols (following a specific language grammar to implement algorithms) leads to the manifestation of complex thought processes. Hence developing code LLMs for both understanding and generation is crucially important both practically and conceptually for stepping towards AGI (Sun et al., 2024a).
Code Foundation Models While many models for code generation are pre-trained mostly on code corpus (Allal et al., 2023; Li et al., 2022b; Fried et al., 2023; Li et al., 2023a), more general purpose LLMs that are continually pre-trained or fine-tuned on code become more powerful and capable such as Codex (Chen et al., 2021b), GPT-4 (OpenAI, 2023a), PaLM-Coder (Chowdhery et al., 2022b), CodeLlama (Rozière et al., 2024), and also smaller scale models like Phi (Gunasekar et al., 2023). The transition from code-specialized to code-understanding models also indicates that coding is a fundamental skill for AGI, just like many other forms of general knowledge. Beyond code generation, these models are also capable of multi-language reasoning (OpenAI, 2023a; Rozière et al., 2024) and infilling with before and after context (Fried et al., 2023; Bavarian et al., 2022). Code models open up many applications as the programs directly serve as the most efficient machine language to communicate with other systems, which we will discuss in the next section.
However, it is worth mentioning that code evaluation is more challenging than pure text for many reasons:
@@
Code LLM Applications Code foundation models (Chen et al., 2021b; OpenAI, 2023a; Rozière et al., 2024; Chowdhery et al., 2022b) have already been extremely capable of conducting many basic code maneuvering such as completion, revision, doc-string generation, commenting, bug finding (Tian et al., 2024b), and code translation (Murali et al., 2024). There are, however, far more exciting applications of these models with no or minimal fine-tuning, which unfolds the possibility of turning many systems into an amalgamation.
@@
  • Software engineering. Many code applications center around software engineering and AI-assisted coding beyond the basic abilities described above. SWE-Bench (Jimenez et al., 2023) attempts to assess a model’s capability to resolve GitHub issues, a core activity in a rich and sustainable real-world software community. Doing so requires a coordinated understanding of the problem description, the execution environment, comments, and the codebase which often has cross-file dependencies and extremely long contexts. The fact that their fine-tuned SWE-Llama can only resolve the simplest issues highly motivates more complicated and capable code models that can greatly help the software ecosystem. Software safety and reliability have always been the most pivotal questions for engineers: RLSQM (Steenhoek et al., 2023) studies using reinforcement learning with static quality metrics as rewards for training a code LLM that can effectively generate unit tests for a codebase with little test smells while adhering to better practices; besides algorithmic bugs,  (Ullah et al., 2023) gives a comprehensive LLM evaluation on identifying security-related bugs, the result of which suggests that even the most capable LLMs like GPT-4 (OpenAI, 2023a) and Palm-2 (Anil et al., 2023) are still prone to non-deterministic responses, incorrect and unfaithful reasoning, and significant non-robustness. LLMs are also explored for lower-level code optimization and refinement. (Cummins et al., 2023) shows that it is possible to fine-tune Llama to optimize LLVM assembly via generating a set of compiler options, which leads to an optimized program and, at the same time, predicts the instruction counts for fine controls. (Wong et al., 2023) investigates the feasibility of utilizing LLMs for de-compiling (reverse-engineer) a C executable into re-compilable C source codes which are expected to exhibit the same functionality, a process that is extremely tedious, time-consuming, and often requires great expertise. Towards a holistic AI development companion, Github Copilot [^17] https://github.com/features/copilot provides personalized and natural language-based coding assistance to developers spanning all levels of expertise and has been integrated into major development workflows. Cognition AI recently announced Devin [^18] https://www.cognition-labs.com/blog as the first AI software engineer. Equipped with its own command line, code editor, and browser, Devin not only achieves the SoTA performance on SWE-bench (Jimenez et al., 2023) but, more impressively, shows its incredible potential in 1) utilizing unfamiliar technology (e.g. running ControlNet on Modal to produce images based on a blog post), 2) building apps end-to-end (e.g. create the Game of Life on a website deployed to Netlify), 3) setting up codes for train and fine-tune LLMs, 4) addressing bugs and feature requests in open source repositories, and so on. Dakhel et al. (2023) suggests examining the capabilities of AI-assisted programming tools in a more controlled setting where the correctness, efficiency, and similarity to human-written solutions are considered extensively.